OpenStudy (-welp-):

|-4+x|-12 What is the value of the expression when x=6

OpenStudy (texaschic101):

| -4 + x | - 12.....when x = 6 sub 6 in for x | -4 + 6 | - 12 = | 2 | - 12 = 2 - 12 = -10

OpenStudy (wil476003):

yo are correct